Solve: 4x/1 + 1 = 6x/6 – 7
Answer by sdmmadam@yahoo.com(530)

Solve: 4x/6 + 1 = 6x/6 – 7
4x/6 + 1 = 6x/6 – 7
1+7 = 6x/6-4x/6(grouping like terms,and when side is changed, sign to bechanged)
8=(6x-4x)/6
8=2x/6
8=x/3
Multiplying by 3
24 = x
Answer: x= 24